OpenStudy (anonymous):

how many real solutions does the equation 36=24x-4x^2 have?

OpenStudy (anonymous):

you want the quick answer?

OpenStudy (anonymous):

or an explanation?

OpenStudy (anonymous):

its my last one so quick

OpenStudy (anonymous):

36=24x-4x^2 4x^2 -24x +36=0 x^2 -6x +9=0 b^2 -4ac = 36-4(9)=0 There is a repeated root

OpenStudy (anonymous):

1 real solution

OpenStudy (anonymous):